# Structure As Search: Unsupervised Permutation Learning for Combinatorial Optimization

### v1

@@ Abstract
We propose a non-autoregressive framework for the Travelling Salesman Problem where solutions emerge directly from learned permutations without explicit search. By applying a similarity transformation to Hamiltonian cycles, the model learns to approximate permutation matrices via continuous relaxations. Our unsupervised approach achieves competitive performance against classical heuristics, demonstrating that the inherent structure of the problem can effectively guide combinatorial optimization without sequential decision-making.
@@ 1 Introduction
How can we solve the Travelling Salesman Problem (TSP), a classic NP-hard problem in combinatorial optimization? Given a set of cities and pairwise distances, the goal is to find the shortest tour that visits each city exactly once and returns to the starting point. A widely held view is that to obtain reasonably good solutions, some form of search is almost always necessary Applegate (2006). Whether through greedy construction, local improvement, or other heuristics, search remains the standard approach. Exact solvers such as Concorde can solve moderate-sized instances optimally, but their exponential runtime limits scalability Applegate et al. (2003). To handle larger instances, local search techniques such as Lin–Kernighan–Helsgaun (LKH) (Lin and Kernighan, 1973; Helsgaun, 2000) have long been used as practical heuristics.
@@
While classical methods remain highly effective, they rely on meticulously designed rules and handcrafted techniques. In contrast, neural network-based approaches aim to learn solution strategies directly from data. One of the earliest neural formulations of TSP was the Hopfield–Tank model (Hopfield and Tank, 1985), which framed the problem as energy minimization in a neural network. Although not data-driven, it marked an early attempt to use neural computation for combinatorial optimization. However, this approach lacked scalability. Recent data-driven methods generally fall into two categories: reinforcement learning (RL), where tours are constructed autoregressively via learned policies (Khalil et al., 2017; Deudon et al., 2018; Kool et al., 2018), and supervised learning (SL), which adopts a two-stage formulation: neural networks generate local preferences, while explicit search procedures build the final solution Joshi et al. (2019). However, RL methods often suffer from sparse rewards and high training variance, while SL approaches require ground-truth solutions during training, which is computationally expensive due to the NP-hard nature of TSP.
@@
In both cases, some form of search is involved, either through learned policies or explicit heuristics. Most RL approaches are also autoregressive, generating tours sequentially in a fixed order. Why might we want to move away from this approach? Many combinatorial optimization problems exhibit strong natural structure. In the case of the TSP, this structure is defined by the search for the shortest Hamiltonian cycle, which inherently constrains the solution space. Recent work (Min et al., 2023) suggests that it is possible to exploit such structure to reduce the search space through an unsupervised, non-autoregressive (NAR) learning. This observation raises a natural question: Is it possible to learn reasonably good TSP solutions without supervision, search, or autoregressive decoding?
@@
In this work, we propose a fundamentally different perspective on learning combinatorial structures. Rather than treating structure as the output of a post-hoc search, we explore the idea that structural inductive bias can replace explicit search . Building on the unsupervised learning for TSP (UTSP) framework (Min et al., 2023; Min and Gomes, 2023), we formulate the TSP as a permutation learning problem: the model directly generates a Hamiltonian cycle using a permutation matrix. Our fully unsupervised, non-autoregressive method requires no optimal training data, search or rollouts. Instead, we train the model using a Gumbel-Sinkhorn relaxation of permutation matrices, followed by hard decoding via the Hungarian algorithm at inference time. This enables solutions to emerge directly from learned structure alone.
@@
We demonstrate that our model consistently outperforms classical baselines, including the nearest neighbor algorithm and farthest insertion, across a range of instance sizes. The structural inductive bias encoded in our model alone generates high-quality solutions without explicit search procedures. Our findings suggest that in combinatorial optimization, appropriately designed structural constraints can serve as effective computational mechanisms, offering a complementary paradigm to conventional search approaches. This indicates that structure itself may be sufficient to guide optimization in certain combinatorial problems.
@@ 2 Background: Unsupervised Learning for the TSP
The key insight is that any Hamiltonian cycle can be generated from the canonical cycle $\mathbb{V}$ through a similarity transformation Min and Gomes (2023). Specifically, if $\mathbf{P}\in S_{n}$ is any permutation matrix, then $\mathbf{P}\mathbb{V}\mathbf{P}^{\top}$ represents a Hamiltonian cycle obtained by reordering the nodes according to permutation $\mathbf{P}$.
Given a distance matrix $\mathbf{D}\in\mathbb{R}^{n\times n}$ where $\mathbf{D}_{ij}$ represents the distance between cities $i$ and $j$, the TSP objective becomes finding the optimal permutation matrix that minimizes:
@@
Since finding the optimal discrete permutation matrix is NP-hard and backpropagating through discrete variables is non-differentiable, we relax the problem by replacing the hard permutation matrix $\mathbf{P}$ with a soft permutation matrix $\mathbb{T}\in\mathbb{R}^{n\times n}$. Following Min et al. (2023); Min and Gomes (2023), we use a Graph Neural Network (GNN) to construct $\mathbb{T}$ and optimize the loss function:
@@ 3 From Soft Permutation $\mathbb{T}$ to Hard Permutation $\mathbf{P}$
To obtain a hard permutation matrix $\mathbf{P}$ from the GNN output, we follow the method proposed in Min and Gomes (2025). We use the Gumbel-Sinkhorn operator Mena et al. (2018), which provides a differentiable approximation to permutation matrices during training. At inference time, we extract a discrete permutation via the Hungarian algorithm. Following the UTSP model Min et al. (2023), the GNN processes geometric node features $f_{0}\in\mathbb{R}^{n\times 2}$ (city coordinates) along with an adjacency matrix $A\in\mathbb{R}^{n\times n}$ defined by:
@@
where $\epsilon$ is i.i.d. Gumbel noise, $\gamma$ is the noise magnitude, $\tau$ is the temperature parameter which controls relaxation sharpness, and $l$ is the number of Sinkhorn iterations. Lower values of $\tau$ yield sharper, near-permutation matrices. At inference time, soft permutation are replaced by a hard permutation via the Hungarian algorithm:
@@ 4.1 Experimental Setup
Our experiments span three distinct problem sizes: 20-node, 50-node, and 100-node TSP instances. For each problem size, we perform hyperparameter sweeps to identify the optimal configurations. Training data consist of uniformly distributed TSP instances generated for each problem size, with 100,000 training instances for both 20-node and 50-node instances, and 300,000 training instances for 100-node instances. All problem sizes use 1,000 instances each for validation and test.
@@ 4.2 Hyperparameter Configuration
We conduct parameter exploration through grid search to evaluate our approach. The 20-node and 50-node instances are trained across all combinations of temperature $\tau\in\{2.0,3.0,4.0,5.0\}$ and noise scale $\gamma\in\{0.005,0.01,0.05,0.1,0.2,0.3\}$, resulting in 24 configurations for each size, while the 100-node instances use only one temperature value $\tau=5.0$ with an expanded noise scale $\gamma\in\{0.005,0.01,0.02,0.03,0.05,0.1,0.2,0.3\}$, totaling 8 configurations. We employ $\ell=60$ Sinkhorn iterations for 20-node instances and $\ell=80$ for 50- and 100-node instances, with training conducted over 300 epochs for 20-node instances and extended to 600 epochs for 50- and 100-node instances to ensure sufficient convergence. For evaluation, tour distances are computed using hard permutations $\mathbf{P}$ obtained via the Hungarian algorithm as described in Equation 9, in contrast to the soft permutation $\mathbb{T}$ used during training.
@@ 4.3 Network Architecture and Training Details
Following the UTSP model (Min et al., 2023), we employ Scattering Attention Graph Neural Networks (SAGs) with 128 hidden dimensions and 2 layers for 20-node instances, 256 hidden dimensions and 4 layers for 50-node instances, and 512 hidden dimensions and 4 layers for 100-node instances (Min et al., 2022). For all instance sizes, SAGs use 6 scattering channels and 2 low-pass channels. We train the networks using the Adam optimizer (Kingma and Ba, 2014) with weight decay regularization $\lambda=1\times 10^{-4}$. Learning rates are set to $1\times 10^{-3}$ for 20-node instances and $2\times 10^{-3}$ for 50- and 100-node instances. To ensure training stability, we implement several regularization techniques: (i) learning rate scheduling with a 15-epoch warmup period, (ii) early stopping with patience of 50 epochs, and (iii) adaptive gradient clipping to maintain stable gradients throughout the optimization process.
@@ 5 Experiment
Our training loss with respect to the validation distance is shown in Figure 1. The training curves consistently converge across all problem sizes, with the training loss (blue) steadily decreasing and stabilizing over epochs. Notably, there is a strong correlation between training loss reduction and validation TSP distance improvement (red), indicating effective learning without overfitting. On 20-node instances, the model achieves the best validation distance of 405.60 at epoch 283; on 50-node instances, our model achieves its best distance of 632.73 at epoch 331; on 100-node instances, we achieve the best validation distance of 874.93 at epoch 519. Across all scales, the validation performance closely tracks the training loss trajectory, confirming that the model generalizes well and that minimizing the objective in Equation 5 consistently leads to improved TSP solution quality.
@@ 5.1 Length Distribution
Figure 3: TSP-50 performance comparison showing our model vs. greedy nearest neighbor baseline. Our model achieves 0.62 shorter tour lengths (mean: 6.37 vs. 6.99) with reduced variability and consistently better performance across all percentiles. Results based on 1,000 test instances.
Figures 2, 3, and 4 show the tour length distributions on the test set for 20-, 50-, and 100-node instances, using the model with the lowest validation length across all hyperparameters. Our model consistently outperforms the Greedy Nearest Neighbor (NN) baseline—which constructs tours by iteratively selecting the nearest unvisited node—achieving substantial gains across all problem sizes. The distribution histograms (left panels) reveal that our model produces more concentratedly distributed, shorter tour lengths with mean values of $\mu=4.06$, $6.37$, and $8.75$ compared to Greedy NN’s $\mu=4.51$, $6.99$, and $9.67$. The box plots (center panels) demonstrate reduced variance and lower median values for our approach, while the cumulative distribution functions (right panels) show our model achieves better solution quality across the entire distribution spectrum, with the curves consistently shifted toward shorter tour lengths.
Figure 4: TSP-100 performance comparison showing our model vs. greedy nearest neighbor baseline. Our model achieves 0.92 shorter tour lengths (mean: 8.75 vs. 9.67) with reduced variability and consistently better performance across all percentiles. Results based on 1,000 test instances.
@@ 5.2 Inference Time
We evaluate the inference efficiency of our approach by measuring the average time per instance, including the GNN forward pass and construction of the hard permutation as defined in Equation 9. On an NVIDIA H100 GPU, the model achieves inference times of 0.19 ms for TSP-20, 0.27 ms for TSP-50, and 0.61 ms for TSP-100.
@@ 5.3 Optimality Gap
[…]
@@
Our method UL, NAR 4.06 6.00% 6.37 12.0% 8.75 12.9%
Our unsupervised, search-free approach demonstrates competitive performance across TSP instances of varying sizes, achieving optimality gaps of 6.00%, 12.0%, and 12.9% on TSP-20, TSP-50, and TSP-100 respectively (Table 1). Notably, our method matches beam search performance on TSP-20 while significantly outperforming it on larger instances (12.0% vs 20.0% gap on TSP-50, and 12.9% vs 27.6% gap on TSP-100). Our approach also consistently outperforms the Greedy NN baseline across all problem sizes, with the performance advantage becoming more pronounced on larger instances.
@@ 6 Hamiltonian Cycle Ensemble
To address this limitation, we revisit our training objective in Equation 5: $\mathcal{L}_{\text{TSP}}=\langle\mathbf{D},\mathbb{T}\mathbb{V}\mathbb{T}^{\top}\rangle$, where our model learns permutations over the canonical Hamiltonian cycle $\mathbb{V}$. We now propose an ensemble approach utilizing powers of the cyclic shift matrix $\mathbb{V}^{k}$, where different values of $k$ satisfying $\gcd(k,n)=1$ generate distinct valid Hamiltonian cycles. Our ensemble strategy trains separate models for each $\mathbb{V}^{k}$ and selects the minimum tour length solution across all $\varphi(n)$ cycle variants for each test instance, where $\varphi(n)$ is Euler’s totient function. This leverages diverse Hamiltonian cycle topologies to mitigate long tail behavior, where although one structure fails, alternatives often succeed, eliminating catastrophic failures.

### v3

@@ Abstract
We propose a non-autoregressive framework for the Travelling Salesman Problem where solutions emerge directly from learned permutations, without requiring explicit search. By applying a similarity transformation to Hamiltonian cycles, the model learns to approximate permutation matrices via continuous relaxations. Our unsupervised approach achieves competitive performance against classical heuristics, demonstrating that the inherent structure of the problem can effectively guide combinatorial optimization without sequential decision-making. Our method offers concrete evidence that neural networks can directly capture and exploit combinatorial structure.
@@ 1 Introduction
How can we solve the Travelling Salesman Problem (TSP), a classic NP-hard problem in combinatorial optimization? Given a set of cities and pairwise distances, the goal is to find the shortest tour that visits each city exactly once and returns to the starting point. A widely held view is that to obtain reasonably good solutions, some form of search is almost always necessary (Applegate, 2006). Whether through greedy construction, local improvement, or other heuristics, search remains the standard approach. Exact solvers such as Concorde can solve moderate-sized instances optimally, but their exponential runtime limits scalability (Applegate et al., 2003). To handle larger instances, local search techniques such as Lin–Kernighan–Helsgaun (LKH) (Lin and Kernighan, 1973; Helsgaun, 2000) have long been used as practical heuristics.
@@
While classical methods remain highly effective, they rely on meticulously designed rules and handcrafted techniques. In contrast, neural network-based approaches aim to learn solution strategies directly from data. One of the earliest neural formulations of TSP was the Hopfield–Tank model (Hopfield and Tank, 1985), which framed the problem as energy minimization in a neural network. Although not data-driven, it marked an early attempt to use neural computation for combinatorial optimization. However, this approach lacked scalability. Recent data-driven methods generally fall into two categories: reinforcement learning (RL), where tours are constructed autoregressively via learned policies (Khalil et al., 2017; Deudon et al., 2018; Kool et al., 2018), and supervised learning (SL), which adopts a two-stage formulation: neural networks generate local preferences, while explicit search procedures build the final solution (Joshi et al., 2019). However, RL methods often suffer from sparse rewards and high training variance, while SL approaches require ground-truth solutions during training, which is computationally expensive due to the NP-hard nature of TSP.
@@
In both RL and SL approaches to the TSP, some form of search is involved, either through learned policies or explicit heuristics. Most RL methods are also autoregressive, generating tours sequentially in a fixed order. Why might we want to move away from this approach? Many combinatorial optimization problems exhibit strong natural structure; in the case of the TSP, this is the shortest Hamiltonian cycle, which inherently constrains the solution space. Recent work (Min et al., 2023) shows that such structure can be exploited through unsupervised, non-autoregressive (NAR) learning, but their framework still requires local search heuristics to assemble final solutions. Building on this, Min and Gomes (2023) formulated TSP as permutation learning with the Gumbel-Sinkhorn operator, though again relying on refinement strategies akin to search. Importantly, Min et al. (2023) demonstrated that unsupervised learning can already reduce the search space, these developments naturally raise the question of whether it is possible to learn high-quality TSP solutions entirely without supervision, search, or autoregressive decoding.
In summary, nearly all data-driven TSP methods, whether supervised, reinforced, or unsupervised, still rely on search. This reliance reveals a core barrier: achieving a learning-based search-free paradigm for TSP remains a major open challenge, as it would provide direct evidence of neural networks’ inherent ability to solve combinatorial problems.
@@
In this work, we propose a different perspective on learning combinatorial structures. Rather than treating structure as the output of a post-hoc search, we explore the idea that structural inductive bias can replace explicit search. Building on the unsupervised learning for TSP (UTSP) framework (Min et al., 2023; Min and Gomes, 2023), we formulate the TSP as a permutation learning problem: the model directly generates a Hamiltonian cycle using a permutation matrix. Our fully unsupervised, non-autoregressive method requires no optimal training data, search or rollouts. Instead, we train the model using a Gumbel-Sinkhorn relaxation of permutation matrices, followed by hard decoding via the Hungarian algorithm at inference time. This enables solutions to emerge directly from learned structure alone.
@@
We demonstrate that our model consistently outperforms classical baselines, including the nearest neighbor algorithm and farthest insertion, across a range of instance sizes. The structural inductive bias encoded in our model alone generates high-quality solutions without explicit search procedures. Our findings suggest that in combinatorial optimization, appropriately designed structural constraints can serve as effective computational mechanisms, offering a complementary paradigm to conventional search approaches. This suggests that structure itself may be sufficient to guide optimization in certain combinatorial problems.
@@ 2 Background: Unsupervised Learning for the TSP
The key insight is that any Hamiltonian cycle can be generated from the canonical cycle $\mathbb{V}$ through a similarity transformation (Min and Gomes, 2023). Specifically, if $\mathbf{P}\in S_{n}$ is any permutation matrix, then $\mathbf{P}\mathbb{V}\mathbf{P}^{\top}$ represents a Hamiltonian cycle obtained by reordering the nodes according to permutation $\mathbf{P}$. Given a distance matrix $\mathbf{D}\in\mathbb{R}^{n\times n}$ where $\mathbf{D}_{ij}$ represents the distance between cities $i$ and $j$, the TSP objective becomes finding the optimal permutation matrix that minimizes:
@@
Since finding the optimal discrete permutation matrix is NP-hard and backpropagating through discrete variables is non-differentiable, we relax the problem by replacing the hard permutation matrix $\mathbf{P}$ with a soft permutation matrix $\mathbb{T}\in\mathbb{R}^{n\times n}$. Following (Min et al., 2023; Min and Gomes, 2023), we use a Graph Neural Network (GNN) to construct $\mathbb{T}$ and optimize the loss function:
@@ 3 From Soft Permutation $\mathbb{T}$ to Hard Permutation $\mathbf{P}$
To obtain a hard permutation matrix $\mathbf{P}$ from the GNN output, we follow the method proposed in (Min and Gomes, 2025). We use the Gumbel-Sinkhorn operator (Mena et al., 2018), which provides a differentiable approximation to permutation matrices during training. At inference time, we extract a discrete permutation via the Hungarian algorithm. Following the UTSP model (Min et al., 2023), the GNN processes geometric node features $f_{0}\in\mathbb{R}^{n\times 2}$ (city coordinates) along with an adjacency matrix $A\in\mathbb{R}^{n\times n}$ defined by:
@@
where $\epsilon$ is i.i.d. Gumbel noise, $\gamma$ is the noise magnitude, $\tau$ is the temperature parameter which controls relaxation sharpness, and $l$ is the number of Sinkhorn iterations. Lower values of $\tau$ yield sharper, near-permutation matrices. At inference, we derive a hard permutation using the Hungarian algorithm:
@@ 4.1 Experimental Setup
Our experiments span three distinct problem sizes: 20-node, 50-node, and 100-node TSP instances. For each problem size, we perform hyperparameter sweeps to identify the optimal configurations. Training data consist of uniformly distributed TSP instances generated for each problem size, with 100,000 training instances for 20-node, 500,000 training instances on 50-node instances, and 1,500,000 training instances for 100-node instances. All problem sizes use 1,000 instances each for validation and test.
@@ 4.2 Hyperparameter Configuration
We conduct parameter exploration through grid search to evaluate our approach. The 20-node instances are trained across all combinations of temperature $\tau\in\{2.0,3.0,4.0,5.0\}$ and noise scale $\gamma\in\{0.005,0.01,0.05,0.1,0.2,0.3\}$, resulting in 24 configurations for each size; the 50-node instances are trained across all combinations of temperature $\tau=5.0$ and noise scale $\gamma\in\{0.005,0.01,0.05,0.1,0.2,0.3\}$, while the 100-node instances use only one temperature value $\tau=5.0$ with an expanded noise scale $\gamma\in\{0.1,0.2,0.3\}$, totaling 3 configurations. We employ $\ell=60$ Sinkhorn iterations for 20-node instances and $\ell=80$ for 50- and 100-node instances, with training conducted over 300 epochs for 20-node instances and extended to 600 epochs for 50- and 100-node instances to ensure sufficient convergence. For evaluation, tour distances are computed using hard permutations $\mathbf{P}$ obtained via the Hungarian algorithm as described in Equation 9, in contrast to the soft permutation $\mathbb{T}$ used during training.
@@ 4.3 Network Architecture and Training Details
Following the UTSP model (Min et al., 2023), we employ Scattering Attention Graph Neural Networks (SAGs) with 128 hidden dimensions and 2 layers for 20-node instances, 256 hidden dimensions and 6 layers for 50-node instances, and 512 hidden dimensions and 8 layers for 100-node instances (Min et al., 2022). For TSP-20, we use SAGs with 6 scattering channels and 2 low-pass channels; for TSP-50 and TSP-100, we use SAGs with 4 scattering channels and 2 low-pass channels. We train the networks using the Adam optimizer (Kingma and Ba, 2014) with weight decay regularization $\lambda=1\times 10^{-4}$. Learning rates are set to $1\times 10^{-3}$ for 20-node instances and $2\times 10^{-3}$ for 50- and 100-node instances. To ensure training stability, we implement several regularization techniques: (i) learning rate scheduling with a 15-epoch warmup period, (ii) early stopping with patience of 50 epochs, and (iii) adaptive gradient clipping to maintain stable gradients throughout the optimization process.
@@ 5 Experiment
Our training loss with respect to the validation distance is shown in Figure 1. The training curves consistently converge across all problem sizes, with the training loss (blue) steadily decreasing and stabilizing over epochs. Notably, there is a strong correlation between training loss reduction and validation TSP distance improvement (red), indicating effective learning without overfitting. On 20-node instances, the model achieves the best validation distance of 405.60 at epoch 283; on 50-node instances, our model achieves its best distance of 627.21 at epoch 550; on 100-node instances, we achieve the best validation distance of 835.38 at epoch 574. Across all scales, the validation performance closely tracks the training loss trajectory, confirming that the model generalizes well and that minimizing the objective in Equation 5 consistently leads to improved TSP solution quality.
@@ 5.1 Length Distribution
Figure 3: TSP-50 performance comparison showing our model vs. greedy nearest neighbor baseline. Our model achieves 0.71 shorter tour lengths (mean: 6.28 vs. 6.99) with reduced variability and consistently better performance across all percentiles. Results based on 1,000 test instances.
Figures 2, 3, and 4 show the tour length distributions on the test set for 20-, 50-, and 100-node instances, using the model with the lowest validation length across all hyperparameters. Our model consistently outperforms the Greedy Nearest Neighbor (NN) baseline—which constructs tours by iteratively selecting the nearest unvisited node—achieving substantial gains across all problem sizes. The distribution histograms (left panels) reveal that our model produces more concentratedly distributed, shorter tour lengths with mean values of $\mu=4.06$, $6.28$, and $8.37$ compared to Greedy NN’s $\mu=4.51$, $6.99$, and $9.67$. The box plots (center panels) demonstrate reduced variance and lower median values for our approach, while the cumulative distribution functions (right panels) show our model achieves better solution quality, with curves consistently shifted toward shorter tour lengths.
Figure 4: TSP-100 performance comparison showing our model vs. greedy nearest neighbor baseline. Our model achieves 1.30 shorter tour lengths (mean: 8.37 vs. 9.67) with reduced variability and consistently better performance across all percentiles. Results based on 1,000 test instances.
@@ 5.2 Inference Time
We evaluate the inference efficiency of our approach by measuring the average time per instance, including the GNN forward pass and construction of the hard permutation as defined in Equation 9. On an NVIDIA H100 GPU with batch size 256, the model achieves inference times of 0.17 ms for TSP-20, 0.15 ms for TSP-50, and 0.40 ms for TSP-100 .
@@ 5.3 Optimality Gap
$\text{Gap (\%)}=\left(\frac{\text{Tour Length (Method)}-\text{Tour Length (Optimal)}}{\text{Tour Length (Optimal)}}\right)\times 100,$ (10)
@@
Our method UL, NAR 4.06 6.00% 6.28 10.4% 8.37 8.00%
Our unsupervised, search-free approach demonstrates competitive performance across TSP instances of varying sizes, achieving optimality gaps of 6.00%, 10.4%, and 8.00% on TSP-20, TSP-50, and TSP-100 respectively (Table 1). Notably, our method matches beam search performance on TSP-20 while significantly outperforming it on larger instances (10.4% vs 20.0% gap on TSP-50, and 8.00% vs 27.6% gap on TSP-100). Our approach also consistently outperforms the Greedy NN baseline across all problem sizes, with the performance advantage becoming more pronounced on larger instances.
@@ 6 Hamiltonian Cycle Ensemble
To address this limitation, we revisit our training objective in Equation 5: $\mathcal{L}_{\text{TSP}}=\langle\mathbf{D},\mathbb{T}\mathbb{V}\mathbb{T}^{\top}\rangle$, where our model learns permutations over the canonical Hamiltonian cycle $\mathbb{V}$. We now propose an ensemble approach utilizing powers of the cyclic shift matrix $\mathbb{V}^{k}$, where different values of $k$ satisfying $\gcd(k,n)=1$ generate distinct valid Hamiltonian cycles. Our ensemble strategy trains separate models for each $\mathbb{V}^{k}$ and selects the minimum tour length solution across all $\varphi(n)$ cycle variants for each test instance, where $\varphi(n)$ is Euler’s totient function. This leverages diverse Hamiltonian cycle topologies to mitigate long tail behavior, so that when one structure fails, alternatives often succeed, thereby eliminating catastrophic failures.